What is the geometric mean of: a and 4a?
What is the geometric mean of: 2a and 4a?

Answer by stanbon(75887)

What is the geometric mean of: a and 4a?
----
Solve for "x":
a/x = x/4a
---
x^2 = 4a^2
x = 2a
--------------------
What is the geometric mean of: 2a and 4a?
Solve for "x":
2a/x = x/4a
x^2 = 8a^2
x = 2a*sqrt(2)
-----------------------
Cheers,
Stan H.
============